# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v2

@@ Abstract
Variable renewable energy droughts, also called “Dunkelflaute”, emerge as a challenge for climate-neutral energy systems based on variable renewables. Drawing on 38 historic weather years and an advanced identification method, we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios. We show that their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts and show how these drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in a perfectly interconnected setting. While the average renewable availability during this period was still 47% of its long-run mean, we argue that system planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using single calendar years is not suitable for modeling weather-resilient energy scenarios.
@@ 1 Introduction
While future energy systems with high VRE shares sources have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflauten”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [raynaud_energy_2018, kittel_measuring_2024] and may cover large geographical areas.If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage and other flexibility options [dowling_role_2020, sepulveda_design_2021]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022, van_der_most_temporally_2024], the U.S. [rinaldi_wind_2021], China [li_renewable_2024], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], Japan [ohba_climatology_2022], and Australia [richardson_climate_2023]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Different methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024, li_renewable_2024]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
@@
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. We draw on a large dataset of historical VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. Our analysis illustrates the variety and seasonality of historical renewable energy drought patterns in Europe and provides evidence on how frequent, how long, and how severe such periods have been in the past, and to what extent they are correlated across the interconnected European energy system in space and time.
@@
Our results quantify how the spatial and temporal complementarity of wind and solar PV alleviates VRE droughts within European countries, giving rise to a technology portfolio effect. Averaged over all thresholds and countries, the maximum drought duration of a renewable technology portfolio that combines solar PV as well as on- and offshore wind power decreases by 64%, 52%, or 47% compared to standalone PV, onshore wind, or offshore wind droughts. Our results also demonstrate how European integration can further mitigate VRE droughts by leveraging spatially complementary VRE availability profiles across regions, signified by a balancing effect. Considering unconstrained geographical balancing across all countries, the longest technology portfolio drought shortens by 65%. We further show that drought characteristics strongly depend on the chosen drought threshold and that single-threshold analyses lead to an incomplete characterization of extreme drought patterns. Using a multi-threshold analysis, we illustrate that the most extreme VRE events manifest as sequential droughts with varying severity, affecting multiple countries simultaneously to different extents. We introduce a drought mass indicator to identify the most extreme droughts and illustrate how these determine long-duration storage needs in a fully renewable European energy system. In Germany, the most extreme event occurred in the winter of 1995/96 and lasted 109 days. Under the assumption of unconstrained geographical balancing across Europe, the most extreme storage-defining event was substantially shorter and lasted 55 days, occurring in the winter of 1996/97. Such extreme droughts may occur at the turn of years, suggesting that planning horizons based on single calendar years are inappropriate for modeling weather-resilient future energy systems.
@@ 2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [kittel_measuring_2024]. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [kittel_measuring_2024]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons (for more details, see Section 4.1). To give an example, an identified wind power drought with a 24-hour duration at a threshold should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
In the main part of the paper, we illustrate selected results for a limited set of countries. The Supplementary Information SI provides additional results. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [kittel_high_2025].
@@ 2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1 shows renewable energy droughts that last longer than one day for the years 1996 and 1997, which emerge to be particularly critical (see Section 2.5). The Figure combines these events for all employed relative thresholds, for selected countries from southern, central and northern Europe, as well as for the pan-European copperplate scenario. These countries represent some of the largest electricity markets in Europe, each with distinct yet complementary wind and solar generation profiles. The pan-European copperplate scenario combines respective profiles from all 33 countries in our data set, assuming unconstrained geographical balancing. The multi-threshold illustration shows the variety of renewable drought patterns in Europe across regions and technologies even within the same time period. It further indicates that the identified drought duration is highly sensitive to the underlying color-coded threshold. In general, single-threshold analyses using low thresholds can detect extreme short-duration droughts, which may occur in isolation or adjacent to high-availability periods (white spaces in the Figure). Higher thresholds increasingly identify longer-lasting events. These potentially include consecutive shorter periods of extreme renewable shortage or surplus, which are smoothed by the averaging mechanics of the identification algorithm (see Section 4.1). The multi-threshold perspective demonstrates that longer-lasting events identified by higher thresholds may encompass multiple severe droughts detected at lower thresholds.
@@
The Figure also illustrates a geographical balancing effect. Assuming perfect interconnection across Europe, geographical balancing can substantially mitigate the severity and duration of renewable energy droughts. This can be seen when comparing the European copperplate scenario (top row in each panel) to the corresponding isolated-countries cases in Figure 1. This balancing effect is particularly pronounced for onshore wind, as onshore wind droughts tend to occur at different times across space, which allows for spatial smoothing across all 33 considered countries. The effect is less pronounced for offshore wind, which can only be deployed in countries with sea access. This limits the spatial extent over which balancing is possible. It is even less pronounced for solar PV because PV droughts are heavily driven by solar seasonality, which affects most countries simultaneously.
Geographical balancing also reduces the severity and duration of technology portfolio drought events as solar seasonality is lower in Southern Europe and wind droughts do not occur simultaneously across countries. Severe and long-lasting portfolio droughts may arise in single countries and also across Europe at the beginning of a calendar year, (e.g., observable in 2012, Figure SI.3), at the end of a year (e.g., 1982, Figure SI.4), or across the turn of years (e.g., in winter 1996/97, Figure 1). In some years, winter portfolio droughts are much less pronounced (e.g., 2013/14, Figure SI.5). Countries relying heavily on wind power, such as the United Kingdom, may also experience very severe events in summer.
While this section highlights important technology-specific and regional characteristics of renewable energy droughts in Europe drawing on specific weather years, it necessarily remains selective. In the following, we analyze the occurrence of European renewable droughts in a more systematic and exhaustive way for all technologies, regions, and weather years in our data, focusing on both shorter and very long-lasting events. In doing so, we also connect to and extend analyses based on single thresholds which have been prevalent in the literature so far.

### v3

# Quantifying the Dunkelflaute — An analysis of variable renewable energy droughts in Europe

@@ Abstract
Variable renewable energy droughts, also called “Dunkelflaute”, emerge as a challenge for climate-neutral energy systems based on variable renewables. Here we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios, using 38 historic weather years and an advanced identification method. Their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, return periods, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts, which drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in an idealized, perfectly interconnected setting. The average renewable availability during this period was still 47% of its long-run mean. System planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using arbitrary single calendar years is not suitable for modeling weather-resilient energy scenarios.
@@ 1 Introduction
While future energy systems with high VRE shares have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflaute”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [raynaud_energy_2018, kittel_measuring_2024] and may cover large geographical areas. If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage as well as other temporal and spatial flexibility options [dowling_role_2020, sepulveda_design_2021]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
@@
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022, van_der_most_temporally_2024, van_duinen_meteorological_2025], the U.S. [rinaldi_wind_2021, bracken_standardized_2024, bracken_seasonal_2025], China [li_renewable_2024, lei_frequency_2024], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], Japan [ohba_climatology_2022], and Australia [richardson_climate_2023]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
@@
Different methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024, li_renewable_2024, biewald_evaluation_2025, wilczak_wind_2025]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. Other researchers have proposed using copulas to construct joint cumulative distribution functions of drought severity and duration [otero_copula-based_2022] or standardized indices to monitor and compare drought events [allen_standardised_2023]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
@@
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. The latter represents an idealized scenario that allows identifying the upper bound of the value of geographical balancing for mitigating renewable energy droughts. We draw on a large dataset of historical VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. Note that the question what mix of storage and other flexibility technologies optimally facilitates weather-resilient renewable energy system is not in the focus of our analysis. This has to be investigated using computationally intensive numerical energy system models. However, the methods and findings presented here can help to identify relevant weather years that should be used in such model analyses.
@@
Our results quantify how the spatial and temporal complementarity of wind and solar PV alleviates VRE droughts within European countries, giving rise to a technology portfolio effect. Averaged over all thresholds and countries, the maximum drought duration of a renewable technology portfolio that combines solar PV as well as on- and offshore wind power decreases by 64%, 52%, or 47% compared to standalone PV, onshore wind, or offshore wind droughts. We also demonstrate how European integration can further mitigate VRE droughts by leveraging spatially complementary VRE availability profiles across regions, signified by a balancing effect. Considering unconstrained geographical balancing across all countries, the longest technology portfolio drought shortens by 65%. We further show that drought characteristics strongly depend on the chosen drought threshold and that single-threshold analyses lead to an incomplete characterization of extreme drought patterns. Using a multi-threshold analysis, we illustrate that the most extreme VRE events manifest as sequential droughts with varying severity, affecting multiple countries simultaneously to different extents. We introduce a drought mass indicator to identify the most extreme droughts and illustrate how these determine long-duration storage needs in a fully renewable European energy system. Considering Germany alone, the most extreme event occurred in the winter of 1995/96 and lasted 109 days. Under the idealized assumption of unconstrained geographical balancing across Europe, the most extreme storage-defining event was substantially shorter and lasted 55 days, occurring in the winter of 1996/97. Such extreme droughts may occur at the turn of years, suggesting that planning horizons based on single calendar years are inappropriate for modeling weather-resilient future energy systems.
@@ 2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [kittel_measuring_2024]. The search algorithm used here iteratively decreases the moving average window, which initially identifies the longest-lasting droughts, followed by progressively shorter ones. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [kittel_measuring_2024]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons. Section 4.1 details the renewable energy drought definition and identification method. To give an example, an identified wind power drought with a 24-hour duration at a threshold should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
In the main part of the paper, we illustrate selected results for a limited set of countries. The SI provides additional results and information on the thresholds used. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [kittel_high_2025].
@@ 2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1 shows renewable energy droughts that last longer than one day for the years 1996 and 1997, which emerge to be particularly critical (see Section 2.6). Figure 1 combines these events for all employed relative thresholds, for selected countries from southern, central and northern Europe, as well as for the pan-European copperplate scenario. These countries represent some of the largest electricity markets in Europe, each with distinct yet complementary wind and solar generation profiles. The pan-European copperplate scenario combines respective profiles from all 33 countries in our data set, assuming unconstrained geographical balancing. Such an idealized scenario is unlikely to be ever achieved in practice, as it would require expanding European grids on a massive scale. However, the copperplate scenario allows identifying the upper bound of the value of European integration to mitigate renewable droughts. The multi-threshold illustration shows the variety of renewable drought patterns in Europe across regions and technologies even within the same time period. It further indicates that the identified drought duration is highly sensitive to the underlying color-coded threshold. In general, single-threshold analyses using low thresholds can detect extreme short-duration droughts, which may occur in isolation or adjacent to high-availability periods (white spaces in the Figure). Higher thresholds increasingly identify longer-lasting events. These potentially include consecutive shorter periods of extreme renewable shortage or surplus, which are smoothed by the averaging mechanics of the identification algorithm (see Section 4.1). The multi-threshold perspective demonstrates that longer-lasting events identified by higher thresholds may encompass multiple severe droughts detected at lower thresholds.
Identified drought patterns in 1996 and 1997 for all investigated relative thresholds with and selected regions. The dates shown on the bottom axis correspond to all technology-region-specific panels. For each panel, the stacked horizontal bands indicate drought occurrences for the color-coded threshold of a particular region. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Figure SI.2 focuses on longer-lasting events by illustrating respective patterns for droughts lasting at least one week.
@@
Figure 1 also illustrates a geographical balancing effect. Assuming perfect interconnection across Europe, geographical balancing can substantially mitigate the severity and duration of renewable energy droughts. This can be seen when comparing the European copperplate scenario (top row in each panel) to the corresponding isolated-countries cases in Figure 1. This balancing effect is particularly pronounced for onshore wind, as onshore wind droughts tend to occur at different times across space, which allows for spatial smoothing across all 33 considered countries. The effect is less pronounced for offshore wind, which can only be deployed in countries with sea access. This limits the spatial extent over which balancing is possible. It is even less pronounced for solar PV because PV droughts are heavily driven by solar seasonality, which affects most countries simultaneously.
Geographical balancing also reduces the severity and duration of technology portfolio drought events as solar seasonality is lower in Southern Europe and wind droughts do not occur simultaneously across countries. Severe and long-lasting portfolio droughts may arise in single countries and also across Europe at the beginning of a calendar year, (e.g., observable in 2012, Figure SI.3), at the end of a year (e.g., 1982, Figure SI.4), or across the turn of years (e.g., in winter 1996/97, Figure 1). This can have important consequences for weather-resilient energy modeling (see Section 3). In some years, winter portfolio droughts are much less pronounced (e.g., 2013/14, Figure SI.5). Countries relying heavily on wind power, such as the United Kingdom, may also experience very severe events in summer.
While this section highlights important technology-specific and regional characteristics of renewable energy droughts in Europe drawing on specific weather years, it necessarily remains selective. Figure SI.6 shows the threshold-specific distribution of all identified droughts across all years in the data. These distributions are highly sensitive to the chosen threshold. While the majority of droughts are relatively short, particularly for lower drought thresholds, higher thresholds find increasingly long-lasting events across all weather years. Further, the distributions confirm the portfolio effect from combining complementary wind and solar PV profiles into a technology portfolio discussed above, both in terms of mean and maximum duration.In the following, we systematically analyze both shorter and very long-lasting renewable droughts for all technologies, regions, and weather years in our data. In doing so, we also connect to and extend analyses based on single thresholds which have been prevalent in the literature so far.